Calculate is the oxidation number of Fe in [Fe(CN)6]4-

Grade:11

4 Answers

Arun
6 years ago

Let the oxidation number of Fe be X

Here oxidation number on CN = -1

For [Fe(CN)6]4- 

X + 6 (-1) = -4

∴ X = +2 

Hence the oxidation state/ number of inner Fe is +2

Rohit
6 years ago
The total charge is -4.The O. State of CN is -1.Therefore the OS of Fe+ OS of(CN)×6 = -4.Therefore OState of Fe is +2.
Ritik Ranjan
6 years ago
Since we consider fe as x.We know that Oxidation number of CN is -1Now, x+(-1)6=-4And solve it .....We get Oxidation state of Fe is +2
Phanindra
6 years ago
Let OS of Fe be x...then....
X + (-1)6 = -4...
X = -4 + 6 …
So x = +2...
So oxidation state of iron Fe in the complex is +2